OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. f(x) = 1(x - 3)^2 + 10

OpenStudy (anonymous):

Medals given

OpenStudy (anonymous):

@zepdrix ?

zepdrix (zepdrix):

Calculate the function values for x=4 and x=7, Take their difference, then divide that difference by the difference in the x values. Wow I said difference a lot +_+

zepdrix (zepdrix):

For linear functions, the rate of change is constant (always stays the same). So this formula might look familiar:\[\Large\rm m=\frac{f(x_2)-f(x_1)}{x_2-x_1}\] For non-linear functions, we call it an average rate of change,\[\Large\rm m_{ave}=\frac{f(7)-f(4)}{7-4}\]It only applies to this specific interval.

OpenStudy (anonymous):

f(4) = 1(4 - 3)^2 + 10 = (4 -3)(4 - 3) + 10 = 12 + 10 = 22 f(7) = 1(7 - 3)^2 + 10 = (7 - 3)(7 - 3) + 10 = 56 + 10 = 66 then its... 66 - 22 / 7 - 4 which is 44 / 3 which is... 14.6666667 ish or 14 2/3??

OpenStudy (anonymous):

@zepdrix ?

zepdrix (zepdrix):

Sorry one sec :)

OpenStudy (anonymous):

its good.

zepdrix (zepdrix):

\[\Large\rm f(4) = 1(\color{red}{4 - 3})^2 + 10\]Do this part in red first. Don't expand out the square first.

OpenStudy (anonymous):

ohhh ok. Lol i see now.

OpenStudy (anonymous):

f(4) = 1 + 10 = 11 and f(7) = 4^2 + 10 = 16 + 10 = 26 so 26 - 11 / 7 - 3 = 15 / 4 = 3 3/4?

zepdrix (zepdrix):

Woops, the denominator is 7-4, right?

OpenStudy (anonymous):

uhh yeah

OpenStudy (anonymous):

oh lol lemme fix that

OpenStudy (anonymous):

so its 15/3 which is 5

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

Thank you so much!